OpenStudy (anonymous):

Solve for y y^2+9y+20=0

OpenStudy (anonymous):

(y+4)(y+5)=0 so y=-4 or -5

OpenStudy (anonymous):

accuzer13 can you explain how you got that answer?

OpenStudy (anonymous):

you factor it to (y+4)(y+5)=0 let y+4=0 so y=-4 and then let y+5=0 so y=-5